# Positivity is undecidable in tensor products of free algebras

### 09 Dec 2023 v1

@@ 4.1. Group stage of construction
[…]
where the first $\mathbb{Z}_{2}$ is generated by $J$, and $\mathbb{Z}_{2}^{\oplus h(m)+1}$ is generated by $\{x_{mi}:i\in\mathbb{Z}_{h(m)+1}\}$. Hence for every $m\in\mathbb{Z}$ there is an automorphism $\phi_{S}^{(m)}:K_{\mathcal{S},x}(m)\rightarrow K_{\mathcal{S},x}(m)$ sending $J\mapsto J$ and $x_{mi}\mapsto x_{m,i+1}$ for all $i\in\mathbb{Z}_{h(m)+1}$, and these extend to the desired automorphism $\phi_{S}:K_{\mathcal{S},x}\rightarrow K_{\mathcal{S},x}$.
The argument for part (b) is identical. For part (c), observe that $K_{\mathcal{S}}(m,0)\cong D_{4}$ for all $m\in\mathbb{Z}$ (since $h(m)\geq 1$), and that $K_{\mathcal{S},0}={}_{\langle J\rangle}\Asterisk_{m\in\mathbb{Z}}K_{\mathcal{S}}(m,0)$ by the normal form theorem for amalgamated free products again. Hence for every $m\in\mathbb{Z}$ there is an isomorphism $\phi_{W}^{(m)}:K_{\mathcal{S}}(m,0)\rightarrow K_{\mathcal{S}}(m+1,0)$ sending $J\mapsto J$, $x_{m,0}\mapsto x_{m+1,0}$, and $z_{m,0}\mapsto z_{m+1,0}$, and these extend to the desired automorphism $\phi_{W}:K_{\mathcal{S},0}\rightarrow K_{\mathcal{S},0}$. ∎
We can now define $G_{\mathcal{S}}$ as the iterated HNN extensions of $K_{\mathcal{S}}$ by the automorphisms $\phi_{S},\phi_{T}$, and $\phi_{W}$ from Lemma 4.5, using variables $S,T$, and $W$ respectively. This means that
$G_{\mathcal{S}}=\langle K_{\mathcal{S}},S,T,W:$ $\ SJS^{-1}=TJT^{-1}=WJW^{-1}=J,$
@@
Unlike $K_{\mathcal{S}}$, this group is finitely generated:
@@
For any $m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}$, we have $x_{mi}=S^{i}W^{m}x_{00}W^{-m}S^{-i}$ and $z_{mi}=T^{i}W^{m}z_{00}W^{-m}T^{-i}$ in $G_{\mathcal{S}}$. ∎
@@
The words $X_{mi}$ and $Z_{mi}$ both have length $12|m|+2|i|+$ 1.
If we restate the presentation of $G_{\mathcal{S}}$ with respect to this generating set, we see that $G_{\mathcal{S}}$ is the group generated by $\mathcal{X}_{G}$, subject to the following relations:
@@
    $[X_{mi},Z_{mi}]=\begin{cases}J&i+1\neq 0\text{ mod }h(m)+1\\ 1&i+1=0\text{ mod }h(m)+1\end{cases}.$
@@
In the next stage of our construction, we want to use Theorem 2.4 to embed $G_{\mathcal{S}}$ into a finitely presented group with polynomial isoperimetric function. For this, we need:
@@
The word problem for $G_{\mathcal{S}}$ with respect to $\mathcal{X}_{G}$ is solvable in deterministic polynomial time.
@@
    If $n=0$, then $w=1$ in $G_{\mathcal{S}}$ if and only if $\eta(g_{0})=1$, and we can stop.
@@
    Otherwise, use Lemma 4.5 to check whether there is an index $1\leq j<n$ such that $e_{j}=-e_{j+1}$, $t_{j}=t_{j+1}=S$, and $g_{j}\in K_{\mathcal{S},x}$. If such a $j$ exists, then let $g_{j}^{\prime}=J^{c}x_{m_{1},i_{1}^{\prime}}\cdots x_{m_{k},i_{k}^{\prime}}$ where $\eta(g_{j})=J^{c}x_{m_{1},i_{1}}\cdots x_{m_{k},i_{k}}$ and $-\lfloor\tfrac{h(m_{l})+3}{2}\rfloor<i_{l}^{\prime}<\lfloor\tfrac{h(m_{l})+2}{2}\rfloor$ is the representative of $[i_{l}+e_{j}]\in\mathbb{Z}_{h(m_{l})+1}$ (computed as in Remark 4.4). Return to step (1) with
@@
    If no such $j$ exists, perform step (4) with $S,K_{\mathcal{S},x}$ replaced by $T,K_{\mathcal{S},z}$ and $W,K_{\mathcal{S},0}$, with $g_{j}^{\prime}$ adjusted to match the automorphisms $\phi_{T}$ and $\phi_{W}$ from Lemma 4.5 respectively.
@@
    If no such $j$ exists for $S,T$ and $W$, then $w\neq 1$ in $G_{\mathcal{S}}$ by Britton’s Lemma, and we can stop.
Define the bit-length and maximum index of a word $w$ over $\widehat{\mathcal{X}}_{G}\cup\widehat{\mathcal{X}}_{G}^{-1}$ analogously to words over $\widetilde{\mathcal{X}}_{K}\cup\widetilde{\mathcal{X}}_{K}^{-1}$. Specifically, if if $w$ is written as in step (1), then define the maximum index of $w$ to be $\max_{i}I_{i}$, where $I_{i}$ is the maximum index of $g_{i}$. Let $N$, $B$, and $I$ be the length, bit-length, and maximum index of the initial input to the above algorithm. Since $n$ decreases by two after each iteration, and $n$ is bounded by $N$, the algorithm is guaranteed to halt after at most $\lceil N/2\rceil$ iterations. In steps (4) and (5), the length of $g_{j}^{\prime}$ is equal to the length of $\eta(g_{j})$, which is at most one more than the length of $g_{j}$. The generators $t_{j}$ and $t_{j+1}$ are removed from $w$ in this step, and $w$ does not change in any other step, so the length of $w$ also decreases in each iteration. The maximum index can increase by one in steps (4) and (5) when the automorphisms $\phi_{S}$ and $\phi_{T}$ are applied, so the maximum index throughout is bounded by $I+\lceil N/2\rceil$. Applying automorphisms $\phi_{S}$, $\phi_{T}$, and $\phi_{W}$ in steps (4) and (5) can also increase the bit-length of each generator by one (note that the bit-length of the representative of $[i]\in\mathbb{Z}_{h(m)+1}$ is at most the bit-length of $i$), so the overall bit length of $w$ can go up by at most $N$ in each step. Thus the bit-length of $w$ is bounded by $B+N\cdot\lceil N/2\rceil$ throughout.
@@
We can now embed $G_{\mathcal{S}}$ into a finitely-presented group:
@@
There is a finitely-presented group $H_{\mathcal{S}}=\langle\mathcal{X}_{H}:\mathcal{R}_{H}\rangle$, a polynomial $d_{\mathcal{S}}$, and an embedding $\iota:\mathcal{F}_{\mathcal{X}_{G}}\hookrightarrow\mathcal{F}_{\mathcal{X}_{H}}$ such that
@@
    $z^{2}\in\mathcal{R}_{H}$ for all $z\in\mathcal{X}_{H}$ (so all the generators of $H_{\mathcal{S}}$ are involutions),
@@
    $\iota(J)$, $\iota(X)$, and $\iota(Z)$ are elements of $\mathcal{X}_{H}$, while $\iota(S)$, $\iota(T)$, and $\iota(W)$ have length two in $H_{\mathcal{S}}$, and
@@
    if $w\in\mathcal{F}_{\mathcal{X}_{G}}$ is a word which is trivial in $G_{\mathcal{S}}$, then there there exist $r_{1},\cdots,r_{k}\in\mathcal{R}_{H}\cup\mathcal{R}_{H}^{-1}$, $z_{1},\cdots,z_{k}\in\mathcal{F}_{\mathcal{X}_{H}}$ with $k\leq d_{\mathcal{S}}(\lvert\iota(w)\rvert)$ and $\lvert z_{i}\rvert\leq d_{\mathcal{S}}(\lvert\iota(w)\rvert)$ for all $1\leq i\leq k$ satisfying
@@
Item (c) implies that $\iota$ embeds $G_{\mathcal{S}}$ into $H_{\mathcal{S}}$.
@@
By Proposition 4.7 and Theorem 2.4, $G_{\mathcal{S}}$ can be embedded in a finitely-presented group $H_{0}=\langle\mathcal{X}_{0}:\mathcal{R}_{0}\rangle$ with polynomial isoperimetric function $d_{H_{0}}$. Let $\mathcal{X}_{1}=\mathcal{X}_{0}\setminus\{J,X,Z\}$, and $\mathcal{X}_{H}=\{J,X,Z\}\cup\{s_{y},t_{y}:y\in\mathcal{X}_{1}\}$. Define $f:\mathcal{F}_{\mathcal{X}_{0}}\to\mathcal{F}_{\mathcal{X}_{H}}$ by $f(y)=y$ for $y\in\{J,X,Z\}$ and $f(y)=s_{y}t_{y}$ for $y\in\mathcal{X}_{1}$, so in other words $f(w)$ is the word $w$ with $y$ replaced by $s_{y}t_{y}$ for all $y\in\mathcal{X}_{1}$. Finally, let $\mathcal{R}_{H}=\{z^{2}:z\in\mathcal{X}_{H}\}\cup\{f(r):r\in\mathcal{R}_{0}\}$. Then
@@
and thus $H_{0}$ is a subgroup of $H:=\langle\mathcal{X}_{H}:\mathcal{R}_{H}\rangle$ by repeatedly applying Lemma 2.1. The inclusion $\iota:G_{0}\hookrightarrow H_{0}\hookrightarrow H$ satisfies parts (a) and (b) of the proposition. Let $\ell$ be the length of the longest relation in $\mathcal{R}_{0}$. If $w\in\mathcal{F}_{\mathcal{X}_{G}}$ is trivial in $G_{\mathcal{S}}$, then $w$ is also trivial in $H_{0}$. Hence by Lemma 2.3, there are $r_{1},\ldots,r_{k}\in\mathcal{R}_{0}\cup\mathcal{R}_{0}^{-1}$ and $z_{1},\ldots,z_{k}\in\mathcal{F}_{\mathcal{X}_{0}}$ such that
@@
where $k\leq d_{0}(|w|)$ and $|z_{i}|\leq\ell k+\ell+|w|$ for all $1\leq i\leq k$. Thus $\iota(w)=f(z_{1})f(r_{1})f(z_{1})^{-1}\cdots f(z_{k})f(r_{k})f(z_{k})^{-1}$. Since $d_{H_{0}}$ is polynomial, $f(r_{i})\in\mathcal{R}_{H}\cup\mathcal{R}_{H}^{-1}$ for all $1\leq i\leq k$, and $|w|\leq|f(w)|\leq 2|w|$ for all $w\in\mathcal{F}_{\mathcal{X}_{0}}$, part (c) is true with e.g. $d_{\mathcal{S}}(n):=2\ell d_{H_{0}}(n)+2\ell+2n$. ∎
4.2. Construction of $\mathcal{A}_{\mathcal{S}}(m)$
For this section, we fix a finitely-presented group $H_{\mathcal{S}}=\langle\mathcal{X}_{H}:\mathcal{R}_{H}\rangle$ and a polynomial $d_{\mathcal{S}}$ as in Proposition 4.8, and identify $G_{\mathcal{S}}$ with its image in $H_{\mathcal{S}}$ (so we don’t mention the embedding $\iota$ explicitly). In particular, if $w\in\mathcal{F}_{\mathcal{X}_{G}}$, then $|w|$ will denote the length of $w$ in $H_{\mathcal{S}}$ rather than $G_{\mathcal{S}}$ unless otherwise noted. The variables $J$, $X$, and $Z$ will continue to denote generators of $H_{\mathcal{S}}$, while $S$, $T$ and $W$ will be length two words in $H_{\mathcal{S}}$. We also continue with the notation $X_{mi}=S^{i}W^{m}XW^{-m}S^{-i}$ and $Z_{mi}=T^{i}W^{m}ZW^{-m}T^{-i}$. Now that we are working with algebras, we let $[A,B]$ denote the algebra commutator $AB-BA$ instead of the group commutator.
To define $\mathcal{A}_{\mathcal{S}}(m)$, consider the set of variables
@@ 4.2. Construction of $\mathcal{A}_{\mathcal{S}}(m)$
in *$\mathbb{C}^{*}\langle\mathcal{X}_{0}\rangle$. For every $m\in\mathbb{N}$, let $\mathcal{A}_{\mathcal{S}}(m)$ be the finitely presented ***-algebra generated by $\mathcal{X}_{H}\cup\mathcal{X}_{0}$, subject to the relations
@@
    $r=1$ for all $r\in\mathcal{R}_{H}$, i.e. all the relations for $H_{\mathcal{S}}$ hold,
@@
    $[\widetilde{X},Q]=\widetilde{X}Q-X_{m,0}Q=0$ and $[\widetilde{Z},Q]=\widetilde{Z}Q-Z_{m,0}Q=0$,
@@
Note that relations (R0) imply that all the generators are order-two unitaries, so $\mathcal{A}_{\mathcal{S}}(m)=\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}_{H}\cup\mathcal{X}_{0}}/\langle\mathcal{R}_{m}\rangle$, where $\mathcal{R}_{m}$ denotes the relations in (R1)-(R6). Since $O_{P}$ and $O_{Q}$ have order two, $P$ and $Q$ are projections in $\mathcal{A}_{\mathcal{S}}(m)$. The point of the generators $O_{P}$ and $O_{Q}$ is to add the projections $P$ and $Q$ while making $\mathcal{A}_{\mathcal{S}}(m)$ a quotient of a free product of $\mathbb{Z}_{2}$’s. Similarly, the point of $U_{1}$ and $U_{2}$ is to add the unitary $U$ without putting any restriction on its order. For the intuition behind the remaining relations, see the representation constructed in the proof of Lemma 4.10 below. Before proceeding to the lemma, we record some facts about the presentation for later use:
@@
If $m\in\mathcal{S}$ then there is a tracial state $\tau$ on $\mathcal{A}_{\mathcal{S}}(m)$ such that $\tau(PQ)>0$.
@@
Suppose $m\in\mathcal{S}$ and let $n:=h(m)<\infty$, so the Turing machine $M$ halts at the $n$-th step on the input $m$. Let $\tau_{0}$ be the canonical trace on $\mathbb{C}H_{\mathcal{S}}$, and let $(\pi:\mathbb{C}H_{\mathcal{S}}\rightarrow\mathcal{B}(\mathcal{H}_{0}),v)$ be the GNS representation of $\tau_{0}$. Let $\mathcal{H}_{1}$ be the Hilbert space $\ell^{2}(\mathbb{Z}_{n+1})$ with the canonical basis $\{e_{i}:i\in\mathbb{Z}_{n+1}\}$, let $E_{i,i}$ be the rank-one projection onto $e_{i}$ for every $i\in\mathbb{Z}_{n+1}$, and let $C:e_{i}\mapsto e_{i+1}$ be the left cyclic shift operator on $\ell^{2}(\mathbb{Z}_{n+1})$. Let ~$\widetilde{\mathcal{H}}=\mathcal{H}_{0}\otimes\mathcal{H}_{1}$, and define a ***-representation $\widetilde{\pi}:\mathbb{C}^{*}\langle\mathcal{X}_{H}\cup\mathcal{X}_{0}\rangle\rightarrow\mathcal{B}(\widetilde{\mathcal{H}}\oplus\widetilde{\mathcal{H}})$ by
@@
$\widetilde{\pi}(U_{1})=\begin{pmatrix}\quad&Id_{\mathcal{H}_{0}}\otimes C\\ Id_{\mathcal{H}_{0}}\otimes C^{-1}\end{pmatrix},\quad\widetilde{\pi}(U_{2})=\begin{pmatrix}\quad&Id_{\widetilde{\mathcal{H}}}\\ Id_{\widetilde{\mathcal{H}}}\end{pmatrix},$
@@
where in the last equation $P_{i}:=(\frac{1-J}{2})(\frac{1-Z_{m,i}}{2})(\frac{1-Z_{m,i+1}}{2})\cdots(\frac{1-Z_{m,n-1}}{2})$ for all $10\leq i\leq n-$ 1, and $P_{n}:=\frac{1-J}{2}$. Note that
$\widetilde{\pi}(U)=\widetilde{\pi}(U_{1})\widetilde{\pi}(U_{2})=\begin{pmatrix}Id_{\mathcal{H}_{0}}\otimes C&\quad\\ \quad&Id_{\mathcal{H}_{0}}\otimes C^{-1}\end{pmatrix}.$
It is not hard to check that $\widetilde{\pi}(r)=0$ for all relations $r$ in (R0)-(R6). Indeed, it is clear that $\widetilde{\pi}(x)$ is a unitary squaring to $1$ for $x\in\mathcal{X}_{H}\cup\{U_{1},U_{2},\widetilde{X},\widetilde{Z}\}$, and $\widetilde{\pi}(Q)$ is a projection. Since $Z_{m,0},\cdots,Z_{m,n-1}$ and $J$ commute in $\mathbb{C}H_{\mathcal{S}}$, $\widetilde{\pi}(P)$ is a projection as well. So $\widetilde{\pi}(r)=0$ for all $r$ in (R0). We also have $\widetilde{\pi}(r)=0$ for $r$ in (R1), (R2), and (R5) by inspection. For (R3),

### 18 Mar 2026 v2

@@ 4.1. Group stage of construction
$\sideset{{}_{\langle J\rangle}}{{}_{i\in\mathbb{Z}_{h(m)+1}}}{\prod}K_{\mathcal{L},x}(m,i)\cong\mathbb{Z}_{2}\times\mathbb{Z}_{2}^{\oplus h(m)+1},$
where the first $\mathbb{Z}_{2}$ is generated by $J$, and $\mathbb{Z}_{2}^{\oplus h(m)+1}$ is generated by $\{x_{mi}:i\in\mathbb{Z}_{h(m)+1}\}$. Hence for every $m\in\mathbb{Z}$ there is an automorphism $\phi_{S}^{(m)}:K_{\mathcal{L},x}(m)\rightarrow K_{\mathcal{L},x}(m)$ sending $J\mapsto J$ and $x_{mi}\mapsto x_{m,i+1}$ for all $i\in\mathbb{Z}_{h(m)+1}$, and these extend to the desired automorphism $\phi_{S}:K_{\mathcal{L},x}\rightarrow K_{\mathcal{L},x}$.
The argument for part (b) is identical. For part (c), observe that $K_{\mathcal{L}}(m,0)\cong D_{4}$ for all $m\in\mathbb{Z}$ (since $h(m)\geq 1$), and that $K_{\mathcal{L},0}={}_{\langle J\rangle}\Asterisk_{m\in\mathbb{Z}}K_{\mathcal{L}}(m,0)$ by the normal form theorem for amalgamated free products again. Hence for every $m\in\mathbb{Z}$ there is an isomorphism $\phi_{W}^{(m)}:K_{\mathcal{L}}(m,0)\rightarrow K_{\mathcal{L}}(m+1,0)$ sending $J\mapsto J$, $x_{m,0}\mapsto x_{m+1,0}$, and $z_{m,0}\mapsto z_{m+1,0}$, and these extend to the desired automorphism $\phi_{W}:K_{\mathcal{L},0}\rightarrow K_{\mathcal{L},0}$. ∎
We can now define $G_{\mathcal{L}}$ as the iterated HNN extensions of $K_{\mathcal{L}}$ by the automorphisms $\phi_{S},\phi_{T}$, and $\phi_{W}$ from Lemma 4.5, using variables $S,T$, and $W$ respectively. This means that
$G_{\mathcal{L}}=\langle K_{\mathcal{L}},S,T,W:$ $\ SJS^{-1}=TJT^{-1}=WJW^{-1}=J,$
@@
Unlike $K_{\mathcal{L}}$, this group is finitely generated:
@@
For any $m\in\mathbb{Z},i\in\mathbb{Z}_{h(m)+1}$, we have $x_{mi}=S^{i}W^{m}x_{00}W^{-m}S^{-i}$ and $z_{mi}=T^{i}W^{m}z_{00}W^{-m}T^{-i}$ in $G_{\mathcal{L}}$. ∎
@@
The words $X_{mi}$ and $Z_{mi}$ both have length $2|m|+2|i|+1$.
If we restate the presentation of $G_{\mathcal{L}}$ with respect to this generating set, we see that $G_{\mathcal{L}}$ is the group generated by $\mathcal{X}_{G}$, subject to the following relations:
@@
    $X_{mi},Z_{mi}]=\begin{cases}J&i+1\neq 0\text{ mod }h(m)+1\\ 1&i+1=0\text{ mod }h(m)+1\end{cases}.$
@@
In the next stage of our construction, we want to use Theorem 2.4 to embed $G_{\mathcal{L}}$ into a finitely presented group with polynomial isoperimetric function. For this, we need:
@@
The word problem for $G_{\mathcal{L}}$ with respect to $\mathcal{X}_{G}$ is solvable in deterministic polynomial time.
@@
    If $n=0$, then $w=1$ in $G_{\mathcal{L}}$ if and only if $\eta(g_{0})=1$, and we can stop.
@@
    Otherwise, use Lemma 4.5 to check whether there is an index $1\leq j<n$ such that $e_{j}=-e_{j+1}$, $t_{j}=t_{j+1}=S$, and $g_{j}\in K_{\mathcal{L},x}$. If such a $j$ exists, then let $g_{j}^{\prime}=J^{c}x_{m_{1},i_{1}^{\prime}}\cdots x_{m_{k},i_{k}^{\prime}}$ where $\eta(g_{j})=J^{c}x_{m_{1},i_{1}}\cdots x_{m_{k},i_{k}}$ and $\lfloor\tfrac{h(m_{l})+3}{2}\rfloor<i_{l}^{\prime}<\lfloor\tfrac{h(m_{l})+2}{2}\rfloor$ is the representative of $i_{l}+e_{j}]\in\mathbb{Z}_{h(m_{l})+1}$ (computed as in Remark 4.4). Return to step (1) with
@@
    If no such $j$ exists, perform step (4) with $S,K_{\mathcal{L},x}$ replaced by $T,K_{\mathcal{L},z}$ and $W,K_{\mathcal{L},0}$, with $g_{j}^{\prime}$ adjusted to match the automorphisms $\phi_{T}$ and $\phi_{W}$ from Lemma 4.5 respectively.
@@
    If no such $j$ exists for $S,T$ and $W$, then $w\neq 1$ in $G_{\mathcal{L}}$ by Britton’s Lemma, and we can stop.
Define the bit-length and maximum index of a word $w$ over $\widehat{\mathcal{X}}_{G}\cup\widehat{\mathcal{X}}_{G}^{-1}$ analogously to words over $\widetilde{\mathcal{X}}_{K}\cup\widetilde{\mathcal{X}}_{K}^{-1}$. Specifically, if if $w$ is written as in step (1), then define the maximum index of $w$ to be $\max_{i}I_{i}$, where $I_{i}$ is the maximum index of $g_{i}$. Let $N$, $B$, and $I$ be the length, bit-length, and maximum index of the initial input to the above algorithm. Since $n$ decreases by two after each iteration, and $n$ is bounded by $N$, the algorithm is guaranteed to halt after at most $\lceil N/2\rceil$ iterations. In steps (4) and (5), the length of $g_{j}^{\prime}$ is equal to the length of $\eta(g_{j})$, which is at most one more than the length of $g_{j}$. The generators $t_{j}$ and $t_{j+1}$ are removed from $w$ in this step, and $w$ does not change in any other step, so the length of $w$ also decreases in each iteration. The maximum index can increase by one in steps (4) and (5) when the automorphisms $\phi_{S}$ and $\phi_{T}$ are applied, so the maximum index throughout is bounded by $I+\lceil N/2\rceil$. Applying automorphisms $\phi_{S}$, $\phi_{T}$, and $\phi_{W}$ in steps (4) and (5) can also increase the bit-length of each generator by one (note that the bit-length of the representative of $i]\in\mathbb{Z}_{h(m)+1}$ is at most the bit-length of $i$), so the overall bit-length of $w$ can go up by at most $N$ in each step. Thus the bit-length of $w$ is bounded by $B+N\cdot\lceil N/2\rceil$ throughout.
@@
We can now embed $G_{\mathcal{L}}$ into a finitely presented group:
@@
There is a finitely presented group $H_{\mathcal{L}}=\langle\mathcal{X}_{H}:\mathcal{R}_{H}\rangle$, a polynomial $d_{\mathcal{L}}$, and an embedding $\iota:\mathcal{F}_{\mathcal{X}_{G}}\hookrightarrow\mathcal{F}_{\mathcal{X}_{H}}$ such that
@@
    $z^{2}\in\mathcal{R}_{H}$ for all $z\in\mathcal{X}_{H}$ (so all the generators of $H_{\mathcal{L}}$ are involutions),
@@
    $\iota(J)$, $\iota(X)$, and $\iota(Z)$ are elements of $\mathcal{X}_{H}$, while $\iota(S)$, $\iota(T)$, and $\iota(W)$ have length two in $H_{\mathcal{L}}$, and
@@
[…]
@@
Item (c) implies that $\iota$ embeds $G_{\mathcal{L}}$ into $H_{\mathcal{L}}$.
@@
By Proposition 4.7 and Theorem 2.4, $G_{\mathcal{L}}$ can be embedded in a finitely presented group $H_{0}=\langle\mathcal{X}_{0}:\mathcal{R}_{0}\rangle$ with polynomial isoperimetric function $d_{H_{0}}$. Let $\mathcal{X}_{1}=\mathcal{X}_{0}\setminus\{J,X,Z\}$, and $\mathcal{X}_{H}=\{J,X,Z\}\cup\{s_{y},t_{y}:y\in\mathcal{X}_{1}\}$. Define $f:\mathcal{F}_{\mathcal{X}_{0}}\to\mathcal{F}_{\mathcal{X}_{H}}$ by $f(y)=y$ for $y\in\{J,X,Z\}$ and $f(y)=s_{y}t_{y}$ for $y\in\mathcal{X}_{1}$, so in other words $f(w)$ is the word $w$ with $y$ replaced by $s_{y}t_{y}$ for all $y\in\mathcal{X}_{1}$. Finally, let $\mathcal{R}_{H}=\{z^{2}:z\in\mathcal{X}_{H}\}\cup\{f(r):r\in\mathcal{R}_{0}\}$. Then
@@
and thus $H_{0}$ is a subgroup of $H:=\langle\mathcal{X}_{H}:\mathcal{R}_{H}\rangle$ by repeatedly applying Lemma 2.1. The inclusion $\iota:G_{0}\hookrightarrow H_{0}\hookrightarrow H$ satisfies parts (a) and (b) of the proposition. Let $\ell$ be the length of the longest relation in $\mathcal{R}_{0}$. If $w\in\mathcal{F}_{\mathcal{X}_{G}}$ is trivial in $G_{\mathcal{L}}$, then $w$ is also trivial in $H_{0}$. Hence by Lemma 2.3, there are $r_{1},\ldots,r_{k}\in\mathcal{R}_{0}\cup\mathcal{R}_{0}^{-1}$ and $z_{1},\ldots,z_{k}\in\mathcal{F}_{\mathcal{X}_{0}}$ such that
@@
where $k\leq d_{0}(|w|)$ and $|z_{i}|\leq\ell k+\ell+|w|$ for all $1\leq i\leq k$. Thus $\iota(w)=f(z_{1})f(r_{1})f(z_{1})^{-1}\cdots f(z_{k})f(r_{k})f(z_{k})^{-1}$. Since $d_{H_{0}}$ is polynomial, $f(r_{i})\in\mathcal{R}_{H}\cup\mathcal{R}_{H}^{-1}$ for all $1\leq i\leq k$, and $|w|\leq|f(w)|\leq 2|w|$ for all $w\in\mathcal{F}_{\mathcal{X}_{0}}$, part (c) is true with e.g. $d_{\mathcal{L}}(n):=2\ell d_{H_{0}}(n)+2\ell+2n$. ∎
4.2. Construction of $\mathcal{A}_{\mathcal{L}}(m)$
For this section, we fix a finitely presented group $H_{\mathcal{L}}=\langle\mathcal{X}_{H}:\mathcal{R}_{H}\rangle$ and a polynomial $d_{\mathcal{L}}$ as in Proposition 4.8, and identify $G_{\mathcal{L}}$ with its image in $H_{\mathcal{L}}$ (so we don’t mention the embedding $\iota$ explicitly). In particular, if $w\in\mathcal{F}_{\mathcal{X}_{G}}$, then $|w|$ will denote the length of $w$ in $H_{\mathcal{L}}$ rather than $G_{\mathcal{L}}$ unless otherwise noted. The variables $J$, $X$, and $Z$ will continue to denote generators of $H_{\mathcal{L}}$, while $S$, $T$ and $W$ will be length two words in $H_{\mathcal{L}}$. We also continue with the notation $X_{mi}=S^{i}W^{m}XW^{-m}S^{-i}$ and $Z_{mi}=T^{i}W^{m}ZW^{-m}T^{-i}$. Now that we are working with algebras, we let $[A,B]$ denote the algebra commutator ABBAAB-BA instead of the group commutator.
To define $\mathcal{A}_{\mathcal{L}}(m)$, consider the set of variables
@@ 4.2. Construction of $\mathcal{A}_{\mathcal{L}}(m)$
in $\mathbb{C}^{*}\langle\mathcal{X}_{0}\rangle$. For every $m\in\mathbb{N}$, let $\mathcal{A}_{\mathcal{L}}(m)$ be the finitely presented *-algebra generated by $\mathcal{X}_{H}\cup\mathcal{X}_{0}$, subject to the relations
@@
    $r=1$ for all $r\in\mathcal{R}_{H}$, i.e. all the relations for $H_{\mathcal{L}}$ hold,
@@
    $\widetilde{X},Q]=\widetilde{X}Q-X_{m,0}Q=0$ and $\widetilde{Z},Q]=\widetilde{Z}Q-Z_{m,0}Q=0$,
@@
Note that relations (R0) imply that all the generators are order-two unitaries, so $\mathcal{A}_{\mathcal{L}}(m)=\mathbb{C}\mathbb{Z}_{2}^{*\mathcal{X}_{H}\cup\mathcal{X}_{0}}/\langle\mathcal{R}_{m}\rangle$, where $\mathcal{R}_{m}$ denotes the relations in (R1)-(R6). Since $O_{P}$ and $O_{Q}$ have order two, $P$ and $Q$ are projections in $\mathcal{A}_{\mathcal{L}}(m)$. The point of the generators $O_{P}$ and $O_{Q}$ is to add the projections $P$ and $Q$ while making $\mathcal{A}_{\mathcal{L}}(m)$ a quotient of a free product of $\mathbb{Z}_{2}$’s. Similarly, the point of $U_{1}$ and $U_{2}$ is to add the unitary $U$ without putting any restriction on its order. For the intuition behind the remaining relations, see the representation constructed in the proof of Lemma 4.10 below. Before proceeding to the lemma, we record some facts about the presentation for later use:
@@
If $m\in\mathcal{L}$ then there is a tracial state $\tau$ on $\mathcal{A}_{\mathcal{L}}(m)$ such that $\tau(PQ)>0$.
@@
[…]
@@
$\widetilde{\pi}(U_{1})=\begin{pmatrix}\quad&Id_{\mathcal{H}_{0}}\otimes\zeta\\ Id_{\mathcal{H}_{0}}\otimes\zeta^{-1}\end{pmatrix},\quad\widetilde{\pi}(U_{2})=\begin{pmatrix}\quad&Id_{\widetilde{\mathcal{H}}}\\ Id_{\widetilde{\mathcal{H}}}\end{pmatrix}$,
@@
where in the last equation $P_{i}:=(\frac{1-J}{2})(\frac{1-Z_{m,i}}{2})(\frac{1-Z_{m,i+1}}{2})\cdots(\frac{1-Z_{m,n-1}}{2})$ for all $0\leq i\leq n-1$, and $P_{n}:=\frac{1-J}{2}$. Note that
$\widetilde{\pi}(U)=\widetilde{\pi}(U_{1})\widetilde{\pi}(U_{2})=\begin{pmatrix}Id_{\mathcal{H}_{0}}\otimes\zeta&\quad\\ \quad&Id_{\mathcal{H}_{0}}\otimes\zeta^{-1}\end{pmatrix}$.
It is not hard to check that $\widetilde{\pi}(r)=0$ for all relations $r$ in (R0)-(R6). Indeed, it is clear that $\widetilde{\pi}(x)$ is a unitary squaring to $1$ for $x\in\mathcal{X}_{H}\cup\{U_{1},U_{2},\widetilde{X},\widetilde{Z}\}$, and $\widetilde{\pi}(Q)$ is a projection. Since $Z_{m,0},\cdots,Z_{m,n-1}$ and $J$ commute in $\mathbb{C}H_{\mathcal{L}}$, $\widetilde{\pi}(P)$ is a projection as well. So $\widetilde{\pi}(r)=0$ for all $r$ in (R0). We also have $\widetilde{\pi}(r)=0$ for $r$ in (R1), (R2), and (R5) by inspection. For (R3),